If I have an 80 in the class and the final exam is worth 15% what is the lowest grade I could get to keep my 80? Isn't it below 80 like can it be a 79?
Found 2 solutions by stanbon, dragonwalker:
Answer by stanbon(75887)

If I have an 80 in the class and the final exam is worth 15% what is the lowest grade I could get to keep my 80?
Your total points for the course are in 2 categores: 85% and 15%
You have 80% of the 85% and you want to know what percent you need of the 15%.
----
Equation:
0.80*0.85 + x%*0.15 = 80%*100%
Multiply thru by 100*100 to get:
80*85 + x*15 = 80*100
15x = 1200
x = 80
=====
You need to get 80% on the final exam.
===================================================
Cheers,
Stan H.
====================

Answer by dragonwalker(73)

If you have an 80 in a class and want to keep it the exact same then you need to get the exact same value no matter how much the exam is worth as a larger score would raise it and a lower score than 80 would lower it. The smaller the value of the exam, the less the change but it would still change if you don't get an exact 80%.
Saying that, if they are rounding up and you actually have a 79.5 or a 80.49 then a very small fluctuation may or may not change the whole number.
To calculate the new score you would need to use the following formula:
current score multiplied by 0.85 + 0.15a = 80, where a is the score of the exam to maintain an 80%.
So 0.15a = 80 - (0.85 x current mark with decimal places if available)
a = (80 - (0.85 x current mark))/0.15
If they round up from 79.5 then you could replace this number with 80 to find the absolute minimum that you will need to maintain an 80 rounded up but I doubt it would make much of a difference. To be safe you should try for an 80.